# Neutron-nucleus dynamics simulations for quantum computers

### 22 Feb 2024 v1

@@ Appendix A Definitions and Lemmas
$k$
$1$
$2$
$4$
$8$
$1$
IIIX
IIXX
IXXX
XXXX
$2$
IIXI
IXXI
XXXI
$3$
IXIX
XXIX
$4$
IXII
XXII
$5$
XIXX
$6$
XIXI
$7$
XIIX
$8$
XIII
@@
$2^{n}+\sum\limits_{k=1}^{K}\overline{n}_{k}2^{n-1},$ (135)
where the first term arises when $k=0$, i.e., the number operators of the form $\{I,Z\}^{\otimes n}$.
@@
$|H(N,K)|=\begin{cases}2^{n}+n2^{n-1}&K=0\\ 2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (136)
@@
Remark 4
For a general potential matrix, the number of terms is given by Lemma A.10. However, the specific structure of the exponential potential $V_{E}(r)$ causes some of the terms to have zero coefficients. In this case, for $K\leq 1$, the diagonal entries of the Hamiltonian are of the form
$H_{i,i}=a+ib,$ (137)
for some constants $a,b$ and for $i\in\{0,\ldots,N-1\}$. Thus, if we consider only the diagonal terms since they are composed of only $\{I,Z\}$, we see that
$H_{\operatorname{diag}}$ $=\sum\limits_{n=0}^{N-1}(a+bn)|n\rangle\!\langle n|$
$\rightarrow\sum\limits_{n=0}^{N-1}(a+nb)O_{e(n)}$
$=\sum\limits_{j=0}^{N-1}\sum\limits_{n=0}^{N-1}(-1)^{e(n)\cdot j}(a+nb)\tilde{O}_{e(n)},$ (138)
where we first map the problem into a particular encoding and the second equality is due to Lemma A.1. In this summation, only terms where $n$ is a power of two survive the summation. Thus, the output consists only $1+\operatorname{log}_{2}(2^{n})=1+n$ terms.
Thus, for the exponential potential, we find that the number of terms is
$|H(N,K)|=\begin{cases}(n+1)+n2^{n-1}&K\in\{0,1\}\\ 2^{n}+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&2\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1}.\end{cases}$ (139)
@@
=$=\begin{cases}2^{n}&K=0\\ 1+\sum\limits_{k=1}^{K}2^{|b(\bar{k})|}\left[1-2^{-\overline{n}_{k}}\right]&1\leq K\leq 2^{n-1}\\ \frac{1}{2}\left(1+3^{n}\right)&K>2^{n-1},\end{cases}$ (145)
@@
Proof. Consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 7 labeled by $2^{j}$. Since the columns represent subsequences ending at index $j$, we see that strings in a column have a fixed structure – the strings end with $XI^{n-j-1}$. The first $j$ bits are a Gray representation of the row index $k-1$. For example, for the cell with row index $3$ and column index $2^{2}$, the entry is of the form
@@
where $U^{\operatorname{GHZ}}$ is the unitary operator
$U^{\operatorname{GHZ}}\coloneqq\prod_{i=|f|}^{2}\operatorname{CNOT}_{1\rightarrow i}H_{1}.$ (163)
The number of two-qubit gates in the diagonalizing unitary for $D(k,n,f)$ is thus given by $(|f|-1)$.
@@
[…]
@@
Proof. As seen in Lemma A.15, the number of two-qubit gates for the diagonalizing unitary for the set $D(n,k,f)$ is $(|f|-1)$. Similar to the proof of Lemma A.13, consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 7 labeled by $2^{j}$. Since the columns represent subsequences ending at $j$, we see that strings in a column have a fixed structure – the strings end with $XI^{n-j-1}$. The first $j$ bits are a Gray representation of the row index $k-1$. For example, the cell with row index $3$ and column index $2^{2}$, the entry is of the form
@@
=$=\begin{cases}0.5n(n-1)&K=0\\ 0.5\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]&1\leq K\leq 2^{n-1}\\ 1+2^{n-1}(n-2)&K>2^{n-1}.\end{cases}$ (181)
Proof. As seen in Lemma A.15, the number of two-qubit gates for the diagonalizing unitary for the set $D(n,k,f)$ is $(|f|-1)$. Similar to the proof of Lemma A.12, Consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 8 labeled by $2^{j}$. For a particular $k$ value, the entry represents a subsequence of length $k$ ending at index $2^{j}$. From the ordering of the binary code basis elements, we see that the entry is given by $b(2^{j+1}-k)$. Thus, the total number is given by
@@ Appendix B List of Operators
Fock operator
Encoded operator
$|0\rangle\!\langle 0|$
$0.5(I-Z_{1})$
$|1\rangle\!\langle 1|$
$0.5(I-Z_{2})$
$|2\rangle\!\langle 2|$
$0.5(I-Z_{3})$
$|3\rangle\!\langle 3|$
$0.5(I-Z_{4})$
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$
$0.5(X_{1}X_{2}+Y_{1}Y_{2})$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$
$0.5(X_{2}X_{3}+Y_{2}Y_{3})$
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$
$0.5(X_{3}X_{4}+Y_{3}Y_{4})$
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$
$0.5(X_{1}X_{3}+Y_{1}Y_{3})$
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$
$0.5(X_{2}X_{4}+Y_{2}Y_{4})$
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$
$0.5(X_{1}X_{4}+Y_{1}Y_{4})$
@@
Fock operator
Encoded operator
$|0\rangle\!\langle 0|$
$0.25(II+IZ+ZI+ZZ)$
$|1\rangle\!\langle 1|$
$0.25(II-IZ+ZI-ZZ)$
$|2\rangle\!\langle 2|$
$0.25(II+IZ-ZI-ZZ)$
$|3\rangle\!\langle 3|$
$0.25(II-IZ-ZI+ZZ)$
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$
$0.5(IX+ZX)$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$
$0.5(XX+YY)$
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$
$0.5(IX-ZX)$
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$
$0.5(XI+XZ)$
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$
$0.5(XI-XZ)$
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$
$0.5(XX-YY)$
@@
Fock operator
Encoded operator
$|0\rangle\!\langle 0|$
$0.25(II+IZ+ZI+ZZ)$
$|1\rangle\!\langle 1|$
$0.25(II+IZ-ZI-ZZ)$
$|2\rangle\!\langle 2|$
$0.25(II-IZ-ZI+ZZ)$
$|3\rangle\!\langle 3|$
$0.25(II-IZ+ZI-ZZ)$
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$
$0.5(XI+XZ)$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$
$0.5(IX-ZX)$
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$
$0.5(XI-XZ)$
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$
$0.5(XX-YY)$
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$
$0.5(XX+YY)$
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$
$0.5(IX+ZX)$
@@ Appendix C Simulation Details for n+C
$N$
Type
Details
n$+^{10}$ $8,16$ Gray Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $500$ iterations, both using SPSA. Lastly, $K=3$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
n$+^{12}$ $8$ Gray Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $500$ iterations, both using SPSA. Lastly, $K=3$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
n$+^{12}$ $16$ Gray Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $1000$ iterations, both using SPSA. Lastly, $K=3$ for $1500$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
n$+^{14}$ $8$ Gray Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $500$ iterations, both using SPSA. Lastly, $K=3$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
n$+^{14}$ $16$ Gray Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $1000$ iterations, both using SPSA. Lastly, $K=3$ for $1500$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $2500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
n$+^{14}$ $8$ OH Noiseless
$K=1$ for $500$ iterations followed by $K=3$ for $1500$ iterations, both using SPSA.
Shot Noise
Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy
$K=1$ for $500$ iterations followed by $K=3$ for $1500$ iterations, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
Table 12: Simulation details for n$+$C. Plots shown in Fig. 11 for n$+^{10}$ Gray, Fig. 12 for n$+^{12}$ Gray, Fig. 13 for n$+^{14}$ Gray, and Fig. 14 for n$+^{14}$ OH.
Appendix D Simulation Details for n$+\alpha$
We now provide specific simulation details for the different simulations in Table 13 for the energy of the lowest $\frac{1}{2}^{+}$ orbit for the n+$\alpha$ optical potential derived ab initio. For all the simulations, we use the Gray encoding, with truncation parameter $N=8$ $(n=3)$ and $N=16$ $(n=4)$, and $K=1,2$. For all Gray encoding simulations, we use $L=5$ layers.
@@ Appendix D Simulation Details for n$+\alpha$
$\hbar\omega$
$N,K$
Type
Details
$12$ $8,1$ Noiseless
@@
$K=0$ for $500$ iterations using SPSA, followed by $K=1$ for $500$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
@@
Same as noiseless but using a shot based estimator for $10^{5}$ shots.
Noisy
@@
$K=0$ for $500$ iterations using $10^{3}$ shots, followed by $K=1$ for $500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$12$ $8,2$ Noiseless
@@
$K=0$ for $500$ iterations, $K=1$ for $500$ iterations, followed by $K=2$ for $500$ iterations, all using SPSA, followed by $K=2$ for $500$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
@@
Same as noiseless but using a shot based estimator for $10^{5}$ shots.
Noisy
@@
$K=0$ for $500$ iterations using $10^{3}$ shots, $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=2$ for $500$ iterations using $10^{5}$ shots, all using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$12$ $16,1$ Noiseless
@@
$K=0$ for $500$ iterations, $K=1$ for $1000$ iterations, both using SPSA, followed by $K=1$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
@@
$K=0$ for $500$ iterations, $K=1$ for $500$ iterations, both using SPSA, following by $K=1$ for $1500$ iterations using gradient descent and a varying learning rate scheme. All estimations use $10^{5}$ shots.
Noisy
@@
$K=0$ for $1000$ iterations using $10^{3}$ shots, followed by $K=1$ for $1500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$12$ $16,2$ Noiseless
@@
$K=0$ for $500$ iterations, $K=1$ for $500$ iterations, $K=2$ for $1000$ iterations, all using SPSA, followed by $K=2$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
@@
$K=0$ for $500$ iterations, $K=1$ for $500$ iterations, $K=2$ for $500$ iterations, all using SPSA, following by $K=2$ for $1500$ iterations using gradient descent and a varying learning rate scheme. All estimations use $10^{5}$ shots.
Noisy
@@
$K=0$ for $500$ iterations using $10^{3}$ shots, $K=1$ for $500$ iterations using $10^{3}$ shots, $K=2$ for $500$ iterations using $10^{3}$ shots, followed by $K=2$ for $500$ iterations using $10^{5}$ shots, all using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$16$ $8,1$ Noiseless
@@
$K=0$ for $500$ iterations using SPSA, followed by $K=1$ for $500$ iterations using gradient descent and a varying learning rate scheme.
Shot Noise
@@
Same as noiseless but using a shot based estimator for $10^{5}$ shots.
Noisy
@@
$K=0$ for $500$ iterations using $10^{3}$ shots, followed by $K=1$ for $500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$16$ $8,2$ Noiseless
$K=0$ for $750$ iterations, $K=1$ for $750$ iterations, followed by $K=2$ for $750$ iterations, all using SPSA.
Shot Noise
@@
Same as noiseless but using a shot based estimator for $10^{5}$ shots.
Noisy
@@
Same as noiseless, but using a fake IBMQ backend ibm_manila with $10^{3}$, $10^{3}$, and $10^{5}$ shots for $K=0,1,2$, respectively, to estimate expectation values. Cost function estimated using a fake IBMQ backend ibm_manila.
Table 13: Simulation details for n$+\alpha$. Plots shown in Fig. 15 for $\hbar\omega=12,N=8$, Fig. 16 for $\hbar\omega=12,N=16$, and Fig. 17 for $\hbar\omega=16,N=8$.

### 15 Mar 2026 v3

@@ 88 IIXX
$k$ 11 22 44 88
11 IIIX IIXX IXXX XXXX
22 IIXI IXXI XXXI
33 IXIX XXIX
44 IXII XXII
55 XIXX
66 XIXI
77 XIIX
88 XIII
@@
$\operatorname{d}(n,K)+\sum\limits_{k=1}^{K}\overline{n}_{k}2^{n-1},$ (165)
where the first term arises when $k=0$, i.e., the number operators of the form $\{I,Z\}^{\otimes n}$ using Lemma A.11.
@@
$|H(N,K)|=\begin{cases}\operatorname{d}(n,1)+n2^{n-1}&K=0\\ \operatorname{d}(n,K)+2^{n-1}\sum\limits_{k=1}^{K}\overline{n}_{k}&1\leq K\leq 2^{n-1}\\ 2^{n-1}(1+2^{n})&K>2^{n-1},\end{cases}$ (166)
@@
$=\begin{cases}2^{n}&K=0\\ 1+\sum\limits_{k=1}^{K}2^{|b(\bar{k})|}\left[1-2^{-\overline{n}_{k}}\right]&1\leq K\leq 2^{n-1}\\ \frac{1}{2}\left(1+3^{n}\right)&K>2^{n-1},\end{cases}$ (172)
@@
Proof. Consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 9 labeled by $2^{j}$. Since the columns represent subsequences ending at index $j$, we see that strings in a column have a fixed structure – the strings end with $XI^{n-j-1}$. The first $j$ bits are a Gray representation of the row index k1k-1. For example, for the cell with row index $3$ and column index $2^{2}$, the entry is of the form
@@
where $U^{\operatorname{GHZ}}_{f}$ is the unitary operator
$U^{\operatorname{GHZ}}_{f}\coloneqq\prod_{i=|f|}^{2}\operatorname{CNOT}_{f_{1}\rightarrow f_{i}}H_{f_{1}}.$ (190)
The number of two-qubit gates in the diagonalizing unitary for $D(k,n,f)$ is thus given by (|f|1)(|f|-1).
@@
$\left\{(\tilde{O}_{l})_{\bar{f}}\otimes X_{f}(\tilde{O}_{m})_{f}:\forall l\in\{0,\ldots,2^{|\bar{f}|}-1\},\ \right.\\ \left.\forall m\in\{0,\ldots,2^{|f|}-1\},\ p(m)\,\mathrm{mod}\,2=0\right\},$ (192)
@@
Proof. As seen in Lemma A.17, the number of two-qubit gates for the diagonalizing unitary for the set $D(n,k,f)$ is (|f|1)(|f|-1). Similar to the proof of Lemma A.15, consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 9 labeled by $2^{j}$. Since the columns represent subsequences ending at $j$, we see that strings in a column have a fixed structure – the strings end with $XI^{n-j-1}$. The first $j$ bits are a Gray representation of the row index k1k-1. For example, the cell with row index $3$ and column index $2^{2}$, the entry is of the form
@@
$=\begin{cases}0.5n(n-1)&K=0\\ 0.5\sum\limits_{k=1}^{K}\overline{n}_{k}\left[2|b(\overline{k})|-1-\overline{n}_{k}\right]&1\leq K\leq 2^{n-1}\\ 1+2^{n-1}(n-2)&K>2^{n-1}.\end{cases}$ (208)
Proof. As seen in Lemma A.17, the number of two-qubit gates for the diagonalizing unitary for the set $D(n,k,f)$ is (|f|1)(|f|-1). Similar to the proof of Lemma A.14, Consider $1\leq K\leq 2^{n-1}$ and consider a column in Table 10 labeled by $2^{j}$. For a particular $k$ value, the entry represents a subsequence of length $k$ ending at index $2^{j}$. From the ordering of the binary code basis elements, we see that the entry is given by $b(2^{j+1}-k)$. Thus, the total number is given by
@@ Appendix B List of Operators
Fock operator Encoded operator
$|0\rangle\!\langle 0|$ $0.5(I-Z_{1})$
$|1\rangle\!\langle 1|$ $0.5(I-Z_{2})$
$|2\rangle\!\langle 2|$ $0.5(I-Z_{3})$
$|3\rangle\!\langle 3|$ $0.5(I-Z_{4})$
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$ $0.5(X_{1}X_{2}+Y_{1}Y_{2})$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$ $0.5(X_{2}X_{3}+Y_{2}Y_{3})$
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$ $0.5(X_{3}X_{4}+Y_{3}Y_{4})$
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$ $0.5(X_{1}X_{3}+Y_{1}Y_{3})$
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$ $0.5(X_{2}X_{4}+Y_{2}Y_{4})$
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$ $0.5(X_{1}X_{4}+Y_{1}Y_{4})$
@@
Fock operator Encoded operator
$|0\rangle\!\langle 0|$ $0.25(II+IZ+ZI+ZZ)$
$|1\rangle\!\langle 1|$ 0.25(IIIZ+ZIZZ)0.25(II-IZ+ZI-$Z$)
$|2\rangle\!\langle 2|$ 0.25(II+IZZIZZ)0.25(II+IZ-ZI-$Z$)
$|3\rangle\!\langle 3|$ 0.25(IIIZZI+ZZ)0.25(II-IZ-ZI+$Z$)
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$ $0.5(IX+ZX)$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$ $0.5(XX+YY)$
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$ 0.5(IXZX)0.5(IX-ZX)
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$ $0.5(XI+XZ)$
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$ 0.5(XIXZ)0.5(XI-XZ)
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$ 0.5(XXYY)0.5($X$-$Y$)
@@
Fock operator Encoded operator
$|0\rangle\!\langle 0|$ $0.25(II+IZ+ZI+ZZ)$
$|1\rangle\!\langle 1|$ 0.25(II+IZZIZZ)0.25(II+IZ-ZI-$Z$)
$|2\rangle\!\langle 2|$ 0.25(IIIZZI+ZZ)0.25(II-IZ-ZI+$Z$)
$|3\rangle\!\langle 3|$ 0.25(IIIZ+ZIZZ)0.25(II-IZ+ZI-$Z$)
$|0\rangle\!\langle 1|+|1\rangle\!\langle 0|$ $0.5(XI+XZ)$
$|1\rangle\!\langle 2|+|2\rangle\!\langle 1|$ 0.5(IXZX)0.5(IX-ZX)
$|2\rangle\!\langle 3|+|3\rangle\!\langle 2|$ 0.5(XIXZ)0.5(XI-XZ)
$|0\rangle\!\langle 2|+|2\rangle\!\langle 0|$ 0.5(XXYY)0.5($X$-$Y$)
$|1\rangle\!\langle 3|+|3\rangle\!\langle 1|$ $0.5(XX+YY)$
$|0\rangle\!\langle 3|+|3\rangle\!\langle 0|$ $0.5(IX+ZX)$
@@ Appendix C Simulation Details for n+C
NN Type Details
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $1500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=3$ for $2500$ iterations using $2\times 10^{4}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
Shot Noise Same as noiseless but using a shot based estimator for $10^{3}$ shots.
Noisy $K=1$ for $500$ iterations followed by $K=3$ for $1500$ iterations, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]
@@
$\hbar\omega$ $N,K$ Type Details
@@
$12$ $8,1$ Noiseless $K=0$ for $500$ iterations using SPSA, followed by $K=1$ for $500$ iterations using gradient descent and a varying learning rate scheme.
@@
Shot Noise Same as noiseless but using a shot based estimator for $10^{5}$ shots.
@@
Noisy $K=0$ for $500$ iterations using $10^{3}$ shots, followed by $K=1$ for $500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
@@
$12$ $8,2$ Noiseless $K=0$ for $500$ iterations, $K=1$ for $500$ iterations, followed by $K=2$ for $500$ iterations, all using SPSA, followed by $K=2$ for $500$ iterations using gradient descent and a varying learning rate scheme.
@@
Shot Noise Same as noiseless but using a shot based estimator for $10^{5}$ shots.
@@
Noisy $K=0$ for $500$ iterations using $10^{3}$ shots, $K=1$ for $500$ iterations using $10^{3}$ shots, followed by $K=2$ for $500$ iterations using $10^{5}$ shots, all using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
@@
$12$ $16,1$ Noiseless $K=0$ for $500$ iterations, $K=1$ for $1000$ iterations, both using SPSA, followed by $K=1$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
@@
Shot Noise $K=0$ for $500$ iterations, $K=1$ for $500$ iterations, both using SPSA, following by $K=1$ for $1500$ iterations using gradient descent and a varying learning rate scheme. All estimations use $10^{5}$ shots.
@@
Noisy $K=0$ for $1000$ iterations using $10^{3}$ shots, followed by $K=1$ for $1500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
@@
$12$ $16,2$ Noiseless $K=0$ for $500$ iterations, $K=1$ for $500$ iterations, $K=2$ for $1000$ iterations, all using SPSA, followed by $K=2$ for $1000$ iterations using gradient descent and a varying learning rate scheme.
@@
Shot Noise $K=0$ for $500$ iterations, $K=1$ for $500$ iterations, $K=2$ for $500$ iterations, all using SPSA, following by $K=2$ for $1500$ iterations using gradient descent and a varying learning rate scheme. All estimations use $10^{5}$ shots.
@@
Noisy $K=0$ for $500$ iterations using $10^{3}$ shots, $K=1$ for $500$ iterations using $10^{3}$ shots, $K=2$ for $500$ iterations using $10^{3}$ shots, followed by $K=2$ for $500$ iterations using $10^{5}$ shots, all using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
@@
$16$ $8,1$ Noiseless $K=0$ for $500$ iterations using SPSA, followed by $K=1$ for $500$ iterations using gradient descent and a varying learning rate scheme.
@@
Shot Noise Same as noiseless but using a shot based estimator for $10^{5}$ shots.
@@
Noisy $K=0$ for $500$ iterations using $10^{3}$ shots, followed by $K=1$ for $500$ iterations using $10^{5}$ shots, both using SPSA. Cost function estimated using a fake IBMQ backend ibm_manila.
$16$ $8,2$ Noiseless $K=0$ for $750$ iterations, $K=1$ for $750$ iterations, followed by $K=2$ for $750$ iterations, all using SPSA.
@@
Shot Noise Same as noiseless but using a shot based estimator for $10^{5}$ shots.
@@
Noisy Same as noiseless, but using a fake IBMQ backend ibm_manila with $10^{3}$, $10^{3}$, and $10^{5}$ shots for $K=0,1,2$, respectively, to estimate expectation values. Cost function estimated using a fake IBMQ backend ibm_manila.
[…]